# GraphEx: A Graph-based Extraction Method for Advertiser Keyphrase Recommendation

### v2

@@ 4.1.1. Traditional Metrics
Ideally, metrics should compare the relevancy of the predictions to the input text without limiting the comparison to a set of predefined labels/keyphrases. However, it is difficult to determine the relevance of predictions without any prior labels. So, while previous research have used human judgement (Yue et al., 2010), we use AI-generated evaluations to evaluate at scale. [^8] The AI predictions were benchmarked against positive buyer judgement and achieved more than 90% alignment, similar to how it was done in (Mishra et al., 2024). We generate prompts for Mixtral 8X7B (Jiang et al., 2024) per item, which contains the item’s title and a set of predicted keyphrases. The response is “yes” or “no” for each keyphrase, indicating whether the keyphrase is relevant to the item or if it is irrelevant.
@@
Once a set of keyphrases is determined as relevant for an item, we filter the keyphrases through the high Search Count threshold. This threshold is determined as the 75th percentile of the descending order of search counts of all unique keyphrases in the category such that 25% of the unique keyphrases are above this limit. The keyphrases whose Search Counts are above the threshold are considered as Relevant Head Keyphrases [^9] High Search Count. otherwise they are considered as Relevant Tail Keyphrases. Note that the head keyphrases determined as irrelevant by the AI are not considered for that item’s evaluation, as even though buyers search the keyphrases in large volume they mostly won’t click on the corresponding item. Hence forth, whenever we use the term head keyphrases it means relevant head keyphrases.
@@
4.2. Performance Results
We compare the models based on the effective (relevant and head) keyphrases that each model recommends. Figure 4 shows the per-model number of keyphrases averaged over all items that are evaluated as relevant or irrelevant by AI, while also distinguishing the head and tail types in the relevant keyphrases. The x-axis shows all the models under comparison. The y-axis in figure 4 shows the average number of keyphrases per item that are irrelevant and relevant head/tail keyphrases, while summing up to the total predictions by each model.
It is evident from figure 4 — as the number of predictions generated by a model increases, the number of irrelevant predictions also tends to rise. The predictions from subsequent models (left to right on the x-axis) are deduplicated against keyphrases already recommended by the previous models. A diverse set of keyphrases is beneficial as it typically results in more engagement, especially if the keyphrases are relevant and are head keyphrases. Based on feedback from sellers, they prefer a balance: they dislike having either too many or too few keywords. They prefer as few irrelevant keywords, while still driving more engagement and having a diverse set of keywords. Thus the ultimate goal for models is to predict a reasonable number of total keyphrases with a higher proportion of relevant head keyphrases and a diverse set of keyphrases. Due to the varying number of predictions by each model, we use one set of metrics to compare the relevant and head keyphrases within each model and another set to compare between different models:
@@ 4.2. Performance Results
Table 2 demonstrates the assessments using both sets of metrics on relevant and head keyphrases. The metrics RRR and RHR are calculated using the GraphEx’s predictions as the denominator. It is important to note that each set of metrics alone don’t offer a comprehensive view. Depending on the variation in total predictions between the two models, the RP and HP tend to favor the model with fewer predictions, while RRR and RHR favor model1 if it has a larger count. We don’t show absolute numbers due to the proprietary nature of data and the models.
@@
For clarity, we first discuss the models that have a much larger number of predictions, as seen in Figure 4 which are SL-emb and fastText. For Table 2, fastText and SL-emb has lower RP and HP (columns $2^{nd}$ and $3^{rd}$) than GraphEx due to their large prediction count. However, we can also see that GraphEx outperforms fastText (except CAT_3) and SL-emb in RRR and RHR (columns $4^{th}$ and $5^{th}$). Thus GraphEx has a lower percentage of irrelevant keyphrases and a higher count of relevant and head keyphrases. CAT_3 is a small metacategory with fewer items and lower buyer interaction, leading to fewer keyphrases. Therefore, creating effective keyphrases for GraphEx becomes difficult and necessitates tailored curation. The models that have much smaller total count of predictions are RE, RE-trank, SL-query, and Graphite. In Table 2, it is evident that these models possess a higher RP compared to GraphEx. This is attributable to their lower number of predictions, which skews the proportions. However, excluding RE and Graphite, the models exhibit a significantly smaller HP than GraphEx. Additionally, these models have much smaller RRR and RHR, as shown by the third and fourth columns (metrics) of the table. Despite Graphite having a slightly higher HP for CAT_3, its RRR and RHR is still lower than that of GraphEx for all categories. Consequently, the models are unlikely to achieve substantial clicks like GraphEx due to the fewer head keyphrases.
fastText SL-emb RE-trank SL-query Graphite RE
CAT_1 1.88x 5.07x 11.6x 8.72x 3.06x 1.57x
CAT_2 2.36x 5.63x 23.9x 12.2x 3.26x 1.57x
CAT_3 1.03x 1.87x 16.5x 3.82x 1.44x 1.11x
Table 3. Relative amount of Diverse (exclusive) Head keyphrases in GraphEx in comparison to other models.
@@
The models RE and RE-trank are simple retrieval techniques, based on recalling the ground truth (item-query combinations with associated buyer activity) with a minimum amount of buyer activity in a short lookback period. The results of RE as seen in Table 2 are mixed, with lower HP in CAT_1, while $2.8\%$ and $15.9\%$ more HP than GraphEx in CAT_2 and CAT_3 respectively. The RRR and RHR of RE is always lower than GraphEx. Albeit their simple nature, both RE and RE-trank are recommenders that are closest to the ground truth in terms of actual buyer-engagement.
@@
While we covered the two aspects of comparison, lower irrelevant and higher head keyphrase counts; diversity is another aspect that determines whether the effective keyphrases generated by a model will bring substantial incremental impact. Therefore, the final metric to compare with other models is the diversity of GraphEx’s predictions. We first separate out (mentioned earlier) the unique or diverse head keywords recommended by each model that are relevant to the item. Table 3 shows the relative amount of diverse keyphrases of each model to the diverse keyphrases of GraphEx (averaged per item). It is evident that GraphEx recommends the highest amount of diverse head keyphrases in constrast to any other model.
4.3. Execution Results
It is important for the models to attain the real-time recommendation and model refresh goals as described in section 1.1. We compare only the XMC models with GraphEx as the REs and SLs (except SL-emb) are simple retrieval techniques implemented in the Spark/Hadoop ecosystem while model inferencing is more complex technique. We examine the models based on Inference Latencies, Model Sizes, and Training times.  [^10] SL-emb inference stages are complex, embedding generation occurs in GPU whereas ANN is done in CPU, thus it is difficult to compare the inference latencies with other models.
Figure 5. Execution performance of fastText, Graphite, and GraphEx. The figure on the left shows Inference latency in milliseconds and the figure on the right shows model sizes in megabytes.
For near real-time recommendation, the Inference Latency of a single input should be in milliseconds. The left image of Figure 5 compares the per-input inference latency of the XMC models and GraphEx. The latencies for each model are computed by amortizing the time taken for prediction over the entire test set. We can see that all the models are within the required limit of 10 ms, but fastText takes more time for a prediction. Graphite and GraphEx’s latencies are comparable for the smaller categories (CAT_2 and CAT_3). The performance of GraphEx is superior, attaining up to 17×17$\times$17 × and up to 13×13$\times$13 × more speed up in contrast to fastText and Graphite on CAT_1. If we infer 20 million items in CAT_1, GraphEx will result in energy savings of 11 hours and 8.5 hours with respect to fastText and Graphite, respectively.
The right of Figure 5 compares the storage sizes of the above models. The fastText model requires significantly more storage across all categories because of the extensive weight matrix and the word embeddings it maintains. This is the case even after reducing the model size during training to enhance precision for production. Graphite occupies substantial space for the large category CAT_1 but has a comparable size to GraphEx for other categories. GraphEx occupies the least size for its models even after constructing graphs for multiple leaf categories. The training times of fastText run into $>4$ hours for all categories with bigger categories running for days and include multiple epochs and autotuning phases. Graphite has a graph-based construction step that takes around $1-6$ minutes while GraphEx takes $<1$ minutes on all the categories. This is due to the curtailment of the training data to head keyphrases and implementation of the construction step that efficiently constructs and stores the model.
Figure 6. GraphEx Batch/NRT Serving Architecture
4.4. Production Engineering Architecture
In this section, we describe the engineering architecture used to serve GraphEx keywords to our sellers for their inventories in one of eBay’s major sites. There are two components for recommendation Batch and Near Real-Time (NRT) Inference. Batch inference primarily serves items with a delay, whereas NRT serves items on an urgent basis, such as items newly created or revised by sellers. The batch inference is done in two parts: 1) for all items in eBay, and 2) daily differential, i.e. the difference of all new items created/revised and then merged with the old existing items. The NRT inference is done using Python code hosted by eBay’s internal ML inference service Darwin. Darwin is then called by eBay’s recommendation service, triggered by the event of new item creation or revision, behind a Flink processing window and feature enrichment. Note that GraphEx serves as one of the keyword recommendation sources in the whole Batch/NRT framework.
The GraphEx batch inference is done using eBay’s machine learning platform Krylov(Katariya and Ramani, 2019) and runs on a single node with 70 cores and 900 GB RAM. GraphEx inference is so fast that the time required to run on a space of 200 million items is just 1.5 hrs. This is a huge improvement over fastText and Graphite which take 1.75 days and 1.5 days respectively. Another batch job in Spark joins these sources in Hadoop and injects them into a Key-Value store (NuKV), which is then called by the eBay platform’s inference api and served to eBay’s sellers. This architecture can scale to billions of items and hundreds of billions of keywords that serve eBay’s platform. The architecture is illustrated in Figure  6. Due to the nature of its algorithm, the GraphEx model is bounded by the label space on which it trains. However, since GraphEx training is inexpensive as Graphite, the model can train in a matter of minutes even for very large categories, making it ideal for daily model refresh. This makes it possible for GraphEx to cater to newer keywords that arise every day. This is a huge improvement over fastText which takes a day or more to train on these large categories and has a monthly refresh schedule.
4.5. Impact
GraphEx was deployed for the sellers of a particular site on eBay to replace Graphite keywords. After its release, a differential pre-post analysis was done to gauge the impact of GraphEx keywords in comparison to Graphite which it replaces. The differential analysis also involved measuring the impact of all keywords generated by GraphEx over a period of 2 weeks, compared to the other sources of recommendations. GraphEx provides 43% more distinct item-keyword associations than Graphite with the average search volume of its keywords nearing 30x of RE, and 2.5X of fastText. In terms of performance, GraphEx delivers an incremental lift of 8.3% in total ads revenue and a 10.3% in Gross Merchandise volume Bought (GMB), i.e. the total money made by selling the item. In terms of Return on Ads Spend (ROAS), given by $ROAS=\frac{GMB}{Ads\;Revenue}$, it is the most successful amongst cold-start models. Among all models its ROAS is only beaten by RE and RE-trank which are non-cold start ground-truth recalling models, and GraphEx beats them in terms of item coverage (more than 3x items covered by GraphEx). We cannot disclose anymore details due to business and proprietary reasons.
5. Conclusion
We introduce a novel graph-based extraction method called GraphEx which is tailored for online advertising in the e-commerce sector. GraphEx efficiently solves the permutation problem of token extraction from item’s title and mapping them to a set of valid keyphrases. It is not limited by the vocabulary of the item’s title and the order of tokens in them. This method produces more item-relevant keyphrases and also targets head keyphrases favored by advertisers, ultimately driving more sales. It is currently implemented at eBay, a leading e-commerce platform serving its sellers with billions of items daily. We show that traditional metrics do not provide accurate comparison amongst the models, and using a single metric for comparison will be misleading. Thus, we use a combination of metrics with AI evaluations to provide a better picture of the practical challenges of keyphrase recommendation. We evaluated its performance against the production models at eBay, demonstrating superior results for our model across the various metrics. Additionally, GraphEx offers the most profitable cold start keyphrase recommendations for advertisers with the lowest inference latency in eBay’s current system and allows for daily model refreshes to serve our ever changing query space.
Appendix A Appendix
A.1. AI Evaluation Prompt
Prompts were generated for Mixtral 8X7B to determine whether a keyphrase was relevant to an item based on similarity to the item’s title as described in Section 4.1.1. The structure of the prompt is shown below. The response is an yes or no answer for each corresponding keyphrase indicating if it is relevant or not relevant to the item.
### Instruction:
Given an item with title: "{title}", determine whether the keyword: "{keyword}", is relevant for cpc targeting or not by giving ONLY yes or no answer:
### Response:
A.2. Ablation Studies
A.2.1. Data Curation Effects
@@
A critical component of GraphEx’s training involves the process of data curation. We find that the Search Count defined in Section 3.2 is crucial for predicting relevant as well as head keyphrases. A low Search Count of 1 inculcates many bogus user queries and hence needs a much higher threshold. An ideal threshold would be keyphrases that are queried at least once daily, which equates to 180 over a span of 6 months. However, as indicated in Table 1, this threshold results in a reduced number of unique keyphrases, necessitating a relaxation of the limit.
@@
To comprehend the influence on recommendations, we evaluated two GraphEx models constructed with search counts of 90 and 180, respectively. A random subset of 1000 items from CAT_1 was utilized for testing. Approximately $20.1\%$ of the items had identical recommendation sets from both models. For the remaining $80\%$ of items, $20\%$ had similar relevant keyphrases and $7.2\%$ had the same relevant head keyphrases. For the remaining keyphrases (about 60%), the proportions of relevant and head keyphrases for the Search Count thresholds of 90 and 180 are presented in Table 4. The benefit obtained with head keyphrases at the 180 search count surpasses the benefit obtained for relevant keyphrases at the same count when compared to a search count of 90.
Search
Count
@@
Table 4. Percentage of relevant and head keyphrases (exclusive) for training curated with different Search Count thresholds.
A.3. Interpretability
The applications in E-Commerce domain frequently require that a model be interpretable. As this helps to comprehend the rational behind its predictions and decision process. In our use case, it is essential to trace where the words in the keyphrases arrive from. Neural Network models typically require converting input text into vectors, which often obscures the contribution of individual tokens to the decisions. While, interpretability techniques such as LIME and SHAP offer post-hoc explanations, treating a Deep Neural Network as a black-box. They also require much effort to figure out the contributions of each input feature.
Unlike the black-box models, the GraphEx algorithm has 3 transparent phases: keyphrase curation, keyphrase mapping, and ranking. The data curation process gives perspective as to how the keyphrases in GraphEx’s label set were curated. The keyphrase mapping phase details how GraphEx’s candidate keyphrases were mapped from the keyphrases extracted from the item’s title to GraphEx’s candidate keyphrases. The ranking algorithm which then ranks the mapped candidates is transparent as well. It uses Label Title Alignment (LTA) outlined in Section 3.5.1 which is a token based algorithm ensuring majority of the tokens in the keyphrases match the title. This ensures that GraphEx’s predicted keyphrases are explainable and interpretable.

### v3

@@ IV-C Evaluation Framework
So, while previous research has used human judgement [12], we use AI-generated evaluations to evaluate at scale for the variety of items at eBay. [^8] The AI predictions were benchmarked against positive buyer judgement and achieved more than 90% alignment, similar to how it was done in [6]. We generate prompts for Mixtral 8X7B [^9] We experimented with GPT4 [38] model and the results are comparable. We use Mixtral because of costs and API rate limits. [39] per item, which contains the item’s title and a set of predicted keyphrases. The structure of the prompt is shown below. The response is “yes” or “no” for each keyphrase, indicating whether the keyphrase is relevant to the item or if it is irrelevant.
### Instruction:
Given an item with title: "{title}", determine whether the keyphrase: "{keyphrase}", is relevant for cpc targeting or not by giving ONLY yes or no answer:
### Response:
@@
Once a set of keyphrases is determined as relevant for an item, we filter the keyphrases through the high Search Count threshold. This threshold is determined as the 90th percentile of the descending order of search counts of all unique keyphrases in the category such that 10% of the unique keyphrases are above this limit. The keyphrases whose Search Counts are above the threshold are considered as Relevant Head Keyphrases [^10] High Search Count. otherwise they are considered as Relevant Tail Keyphrases. [^11] This evaluation framework is only for offline analysis and none of the model’s recommendations to the seller are filtered in this way as it would be infeasible to do so due to the latency of LLMs. One intuition behind this evaluation is that head keyphrase determined as irrelevant by the AI will be ineffective — even though buyers search the keyphrases in large volume, they mostly will not click on the corresponding item.
We compare the models based on the effective (relevant and head) keyphrases that each model recommends. Figure 4 shows the per-model number of keyphrases averaged over all items that are evaluated as relevant or irrelevant by AI, while also distinguishing the head and tail types in the relevant keyphrases. The x-axis shows all the models under comparison. The y-axis shows the average number of keyphrases per item that are irrelevant and relevant head/tail keyphrases, while summing up to the total predictions by each model.
@@
It is evident from Figure 4 — as the number of predictions generated by a model increases, the number of irrelevant predictions also tends to rise. Our evaluation metrics are as follows:
@@
Due to the varying number of predictions by each model, we use one set of metrics to compare the relevant and head keyphrases within each model (RP and HP) and another set to compare between different models (RRR and RHR).
IV-D Performance Results
We use the metrics defined in the previous section to compare each model’s predictions.
IV-D1 Performance Comparison
@@
Table III demonstrates the evaluations using both sets of metrics on relevant and head keyphrases. The metrics RRR and RHR are calculated using the GraphEx’s predictions as the denominator (model2). It is important to note that each set of metrics alone do not offer a comprehensive view. Depending on the variation in the total predictions between the two models, RP and HP tend to favor the model with fewer predictions, while RRR and RHR favor model1 if it has a higher count. We do not show absolute numbers due to the proprietary nature of data and the models.
@@
For clarity, we first discuss the models that have a much large number of predictions, as seen in Figure 4 which are SL-emb and fastText. For Table III, fastText and SL-emb have lower RP and HP (columns $2^{nd}$ and $3^{rd}$) than GraphEx due to their large prediction count. However, we can also see that in RRR and RHR (columns $4^{th}$ and $5^{th}$), GraphEx outperforms fastText (except CAT_3) and SL-emb in all categories. Thus, GraphEx has a lower percentage of irrelevant keyphrases and a higher count of relevant and head keyphrases. CAT_3 is a small metacategory with fewer items and lower buyer interaction, leading to fewer keyphrases. Therefore, creating effective keyphrases for GraphEx becomes difficult and necessitates tailored curation.
The models that have a comparatively smaller total count of predictions are RE, SL-query, and Graphite. In Table III, it is evident that these models (except SL-query) have a higher RP compared to GraphEx. This is attributable to their lower number of predictions, which skews the proportions. However, excluding RE and Graphite, all models exhibit HP significantly smaller than that of GraphEx. Additionally, all the models have much smaller RRR and RHR. Although Graphite has a slightly higher HP for CAT_3, its RRR and RHR are still lower than that of GraphEx for all categories. Consequently, models are unlikely to achieve substantial clicks like GraphEx due to the fewer head keyphrases.
@@
The model RE is a simple retrieval technique, based on recalling the ground truth (item-query combinations with associated buyer activity) with a minimum amount of buyer activity in a short lookback period. The results of RE as seen in Table III are mixed, with lower HP in CAT_1, while $2.8\%$ and $15.9\%$ more HP than GraphEx in CAT_2 and CAT_3 respectively. The RRR and RHR of RE is always lower than GraphEx. Albeit its simple nature, RE is a 100% recall recommender that reflects the ground truth in terms of actual buyer-engagement.
IV-D2 Diversity Comparison
While we covered the two aspects of comparison, lower irrelevant and higher head keyphrase counts; diversity is another aspect that determines whether the effective keyphrases generated by a model will bring substantial incremental impact. In complicated systems like e-commerce where we have multiple retrieval sources the incremental impact of any model is dictated by the amount of clicks/sales/revenue brought about by only the keyphrases that are unique to the model, i.e., not present in any other retrieval sources. A diverse set of keyphrases is beneficial as it typically results in more engagement, especially if the keyphrases are relevant and are head keyphrases. Based on feedback from sellers, they prefer a balance: they dislike having either too many or too few keyphrases. Thus the ultimate goal is to predict a reasonable number of total keyphrases with a higher proportion of relevant keyphrases and a diverse set of head keyphrases.
Figure 5: Venn diagram representing the different recall sources and their keyphrases. The intersecting regions depect the keyphrases from multiple recall sources, while the non-intersecting crossed out regions depict the keyphrases belonging solely to that recall source — which contribute to the incremental impact which each recall source brings.
@@ IV-D2 Diversity Comparison
Therefore, the final metric to compare with other models is the exclusive diversity of GraphEx’s predictions. For this evaluation, as illustrated in Figure 5, we first separate out the unique head keyphrases recommended by each model that are relevant to the item. Table IV shows the relative amount of relevant exclusive head keyphrases of each model to the relevant exclusive head keyphrases of GraphEx (averaged per item). It is evident that GraphEx recommends the highest amount of diverse relevant head keyphrases in contrast to any other model.
Models CAT_1 CAT_2 CAT_3
fastText 1.88x 2.36x 1.03x
SL-emb 5.07x 5.63x 1.87x
SL-query 8.72x 12.2x 3.82x
Graphite 3.06x 3.26x 1.44x
RE 1.57x 1.57x 1.11x
TABLE IV: Relative amount of relevant diverse (exclusive) Head keyphrases in GraphEx in comparison to other models. Each model’s predictions are standalone.
IV-E Miscellaneous Results
For sake of completeness, we compare the precision and recall scores of all the models. As discussed earlier, RE uses the click-based keyphrases for each item, hence we use RE’s recommended keyphrases as the ground truths to compare other models’ predictions with it. Table V shows the precision and recall scores of each model relative to GraphEx.
Metrics fastText Graphite Sl-emb SL-query
Precision 1.08 1.84 0.87 0.95
Recall 1.09 1.62 4.01 3.43
TABLE V: Relative Precision and Recall numbers of other models with respect to Graphex.
The numbers reflect that GraphEx has the lowest recall and low precision scores, indicating minimal ground truth retrieval capability. This works in its favor as the de-duplication with RE (recall) is minimal. This is further reinforced in the impact where GraphEx’s incremental impact is superior to the other models.
IV-F Abalation Studies
IV-F1 Alignment functions
We experimented with various alignment functions other than LTA, such as Word Match Ratio (WMR) used by Graphite [6] and the Jaccard coefficient (JAC). In terms of the notations used for LTA in section III-E1, we re-define as $WMR=\frac{c}{|l|}$ and $JAC=\frac{c}{|l|+|T|-c}$. Table VI compares the results when we use the three alignment functions in the GraphEx algorithm and generate recommendations for the sample of items in the previous section.
It might seem like JAC is the same as LTA, since the only differentiation is the term $|T|$ in the denominator of JAC. This term is constant when re-ranking the candidate keyphrases for a single input title. Hence both LTA and JAC should behave similarly, which is somewhat true from table VI. However, there is still a difference because $|T|$ is much larger than $|l|$ which generally makes the numerator differentiate among the keyphrases. Let’s take an example title with 10 tokens (A-J) which we will compare with two keyphrases “A B C” and “A B C D E”, LTA ranks higher the former ($\frac{3}{1}>\frac{4}{2}$) while JAC ranks the latter higher ($\frac{3}{10}<\frac{4}{10}$). The token “E” is risky, as it can be incorrect or change the product entirely; this risk is minimized by using LTA which penalizes such mismatches. Large-scale experiments with over 110 million items have shown at least 5% points difference in RP between both the functions.
Category RP (%)
WMR JAC LTA
CAT_1 33.6 44.5 45.8
CAT_2 40.8 40.8 40.8
CAT_3 42.6 55.0 56.0
TABLE VI: Table showing the relevant proportion (RP) of Word Match Ratio (WMR), Jaccard Coefficient(JAC) and Label Title Alignment (LTA) when used in GraphEx algorithm for the 3 categories.
Additional alignment techniques such as semantic matching were considered but avoided due to latency constraints. We experimented with subword matching to improve our token matching function; this increased the inference latency without too much improvement in performance. We used a proprietary stemming function for words to increase the reach of token matches.
IV-F2 Data Curation Effects
@@
A critical component of GraphEx’s training involves the process of data curation. We find that the Search Count defined in Section III-B is crucial for predicting relevant as well as head keyphrases. A low Search Count of 1 inculcates many bogus user queries and hence needs a much higher threshold. An ideal threshold would be keyphrases that are queried at least once daily, which equates to 180 over a span of 6 months. However, as indicated in Table II, this threshold results in a reduced number of unique keyphrases, necessitating a relaxation of the limit.
@@
To comprehend the influence on recommendations, we evaluated two GraphEx models constructed with search counts of 90 and 180, respectively. A random subset of 1000 items from CAT_1 was utilized for testing. Approximately $20.1\%$ of the items had identical recommendation sets from both models. For the remaining $80\%$ of items, $20\%$ had similar relevant keyphrases and $7.2\%$ had the same relevant head keyphrases. For the remaining disparate recommendations (about 60%), the proportions of relevant and head keyphrases for the Search Count thresholds of 90 and 180 are presented in Table VII. The benefit obtained with head keyphrases at the 180 search count surpasses the benefit obtained for relevant keyphrases at the same count when compared to a search count of 90.
Search Count
@@
TABLE VII: Percentage of relevant and head keyphrases (exclusive) for training curated with different Search Count thresholds.
IV-G Execution Results
It is important for the models to attain the real-time recommendation and model refresh goals as described in section I-A5. We compare only the XMC models with GraphEx as the REs and SLs (except SL-emb) are simple retrieval techniques implemented in the Spark/Hadoop ecosystem while model inferencing is more complex technique. We examine the models based on Inference Latencies, Model Sizes, and Training times.  [^12] SL-emb inference stages are complex, embedding generation occurs in GPU whereas ANN is done in CPU, thus it is difficult to compare the inference latencies with other models.
(a) Inference latency in milliseconds.
(b) Model sizes in megabytes.
Figure 6: Execution performances of fastText, Graphite, and GraphEx.
For near real-time recommendation, the Inference Latency of a single input should be in milliseconds. The Figure 6(a) compares the per-input inference latency of the XMC models and GraphEx. The latencies for each model are computed by amortizing the time taken for prediction over the entire test set. We can see that all the models are within the required limit of 10 ms, but fastText takes more time for a prediction. Graphite and GraphEx’s latencies are comparable for the smaller categories (CAT_2 and CAT_3). The performance of GraphEx is superior, attaining up to 17×17$\times$17 × and up to 13×13$\times$13 × more speed up in contrast to fastText and Graphite on CAT_1. If we infer 20 million items in CAT_1, GraphEx will result in energy savings of 11 hours and 8.5 hours with respect to fastText and Graphite, respectively.
The Figure 6(b) compares the storage sizes of the above models. The fastText model requires significantly more storage across all categories because of the extensive weight matrix and the word embeddings it maintains. This is the case even after reducing the model size during training to enhance precision for production. Graphite occupies substantial space for the large category CAT_1 but has a comparable size to GraphEx for other categories. GraphEx occupies the least size for its models even after constructing graphs for multiple leaf categories. The training times of fastText run into $>4$ hours for all categories with bigger categories running for days and include multiple epochs and autotuning phases. Graphite has a graph-based construction step that takes around $1-6$ minutes while GraphEx takes $<1$ minutes on all the categories. This is due to the curtailment of the training data to head keyphrases and implementation of the construction step that efficiently constructs and stores the model.
IV-H Production Engineering Architecture
In this section, we describe the engineering architecture used to serve GraphEx keyphrases to our sellers for their inventories in one of eBay’s major sites. There are two components for recommendation Batch and Near Real-Time (NRT) Inference. Batch inference primarily serves items with a delay, whereas NRT serves items on an urgent basis, such as items newly created or revised by sellers. The batch inference is done in two parts: 1) for all items in eBay, and 2) daily differential, i.e. the difference of all new items created/revised and then merged with the old existing items. The NRT inference is done using Python code hosted by eBay’s internal ML inference service Darwin. Darwin is then called by eBay’s recommendation service, triggered by the event of new item creation or revision, behind a Flink processing window and feature enrichment. Note that GraphEx serves as one of the keyphrase recommendation sources in the whole Batch/NRT framework.
Figure 7: GraphEx Batch/NRT Serving Architecture
GraphEx’s batch inference is done using eBay’s machine learning platform Krylov[40] and runs on a single node with 70 cores and 900 GB RAM. GraphEx inference is so fast that the time required to run on a space of 200 million items is just 1.5 hrs. This is a huge improvement over fastText and Graphite which take 1.75 days and 1.5 days respectively. Another batch job in Spark joins these sources in Hadoop and injects them into a Key-Value store (NuKV), which is then called by the eBay platform’s inference api and served to eBay’s sellers. This architecture can scale to billions of items and hundreds of billions of keyphrases that serve eBay’s platform. The architecture is illustrated in Figure  7. Due to the nature of its algorithm, the GraphEx model is bounded by the label space on which it trains. However, since GraphEx training is as inexpensive as Graphite, the model can train in a matter of minutes even for very large categories, making it ideal for daily model refresh. This makes it possible for GraphEx to cater to newer keyphrases that arise every day. This is a huge improvement over fastText which takes a day or more to train on these large categories and has a monthly refresh schedule.
IV-I Impact
GraphEx was deployed for the sellers of a particular site on eBay to replace Graphite keyphrases. After its release, a differential pre-post analysis was done to gauge the impact of GraphEx keyphrases in comparison to Graphite which it replaces. The differential analysis also involved measuring the impact of all keyphrases generated by GraphEx over a period of 2 weeks, compared to the other sources of recommendations. GraphEx provides 43% more distinct item-keyphrase associations than Graphite with the average search volume of its keyphrases nearing 30x of RE, and 2.5X of fastText. In terms of performance, GraphEx delivers an incremental lift of 8.3% in total ads revenue and a 10.3% in Gross Merchandise volume Bought (GMB), i.e. the total money made by selling the item. In terms of Return on Ads Spend (ROAS), given by $ROAS=\frac{GMB}{Ads\;Revenue}$, it is the most successful among the cold-start models. Among all models its ROAS is only beaten by RE which are non-cold start 100% recall models, and GraphEx beats them in terms of item coverage (more than 3x items covered by GraphEx). We cannot disclose anymore details due to business and proprietary reasons.
V Conclusion
We introduce a novel graph-based extraction method called GraphEx which is tailored for online advertising in the e-commerce sector. GraphEx efficiently solves the permutation problem of token extraction from item’s title and mapping them to a set of valid keyphrases. It is not limited by the vocabulary of the item’s title and the order of tokens in them. This method produces more item-relevant keyphrases and also targets head keyphrases favored by advertisers, ultimately driving more sales. It is currently implemented at eBay, a leading e-commerce platform serving its sellers with billions of items daily. We show that traditional metrics do not provide accurate comparison amongst the models, and using a single metric for comparison will be misleading. Thus, we use a combination of metrics with AI evaluations to provide a better picture of the practical challenges of keyphrase recommendation. We evaluated its performance against the production models on eBay, demonstrating superior results for our model across the various metrics. Additionally, GraphEx offers the most profitable cold start keyphrase recommendations for advertisers with the lowest inference latency in eBay’s current system and allows for daily model refreshes to serve our ever-changing query space.